7-9 YIELD TO MATURITY Harrimon Industries bonds have 6 years left to maturity. Interest is paid annually, and the bonds have a $1,000 par value and a coupon rate of 10%. a. What is the yield to maturity at a current market price of (1) $865 and (2) $1,166? b. Would you pay $865 for each bond if you thought that a “fair” market interest rate for such bonds was 12%—that is, if rd 12%? Explain your answer.

PLEASE SHOW HOW TO COMPLETE IN EXCEL
Answer #1

Given for Harrimon Industries' bond,

Face value = $1000

coupon rate = 10%

Coupon = 10% of 1000 = $100

years to maturity = 6

a). 1). When current market price is $865

using financial calculator to compute YTM. Use values

FV = 1000

PV = -865

PMT = 100

N = 6

compute for I/Y, we get I/Y = 13.42

So, Yield to maturity at price of $865 is 13.42%.

2). When current market price is $1166

using financial calculator to compute YTM. Use values

FV = 1000

PV = -1166

PMT = 100

N = 6

compute for I/Y, we get I/Y = 6.56

So, Yield to maturity at price of $1166 is 6.56%.

b). when interest rate = 12%,

Calculating price of bond using financial calculator, use values:

FV = 1000

I/Y = 12

PMT = 100

N = 6

compute for PV, we get PV = 917.77

fair price of bond should be 917.77

but since bond is selling at a lower price, it is beneficial to buy the bond. Also, if this bond is held till maturity, actual return of this is 13.42% which is more than the current interest rate of similar bonds.